OpenStudy (anonymous):

can someone tell me how to do this? Find all solutions in the interval [0, 2π). 2 sin2x = sin x

OpenStudy (anonymous):

\[ 4 \sin(x) \cos(x)=\sin(x) \\ 4 \sin(x) \cos(x)-\sin(x)=0 \\ \sin(x) (4 \cos(x) -1)=0\\ \sin(x) =0 \\ 4 \cos(x) -1 =0\\ \cos(x) =\frac 1 4 \] Can you finish it now?

OpenStudy (anonymous):

still confused >.<

OpenStudy (anonymous):

When is sin(x) =0 for x between 0 and 2 Pi?